OpenStudy (iwanttogotostanford):

PLEASE PLEASE HELP WITH THIS QUESTION:

OpenStudy (iwanttogotostanford):

@brainzonly @Kevin @inkyvoyd

OpenStudy (3mar):

May I help?

OpenStudy (kevin):

please

OpenStudy (3mar):

Direct substitution means that replace every x in the limit with its value.

OpenStudy (3mar):

Do a try for the first one.

OpenStudy (iwanttogotostanford):

@Kevin

OpenStudy (iwanttogotostanford):

@Nnesha

OpenStudy (iwanttogotostanford):

someone please help

OpenStudy (3mar):

Let's do it together

OpenStudy (kevin):

just plug x value into the equation. If the result is not equal to '0', it's the answer. IF the result is equal to '0', then you need to factorize the equation or simplify it.

OpenStudy (kevin):

do u understand?

OpenStudy (iwanttogotostanford):

not really no

OpenStudy (iwanttogotostanford):

I'm more confused on these questions than the last

OpenStudy (iwanttogotostanford):

I think the first one is 7 and the second is 1.... am I right??? @Kevin

OpenStudy (iwanttogotostanford):

@mathmate @welshfella plz help

OpenStudy (3mar):

@Kevin

OpenStudy (kevin):

the second is not 1 can u give me know how you get 1?

OpenStudy (kevin):

@iwanttogotostanford

OpenStudy (brainzonly):

ummmm.....I don't see the question?..

OpenStudy (iwanttogotostanford):

it is attached as a screenshot a couple above

OpenStudy (iwanttogotostanford):

i thought the first was 7 and the second was 1

OpenStudy (kevin):

the second is not 1 can you write your calculation here

OpenStudy (kevin):

\[\frac{ x^2 - 4 }{ x-2 }\] If you fill x with 2, then it would be : 0/0 That's mean you need to factorize it : \[\frac{ (x + 2)(x-2) }{ x-2 }\] you can divide x-2 by x-2, the quotient would be: x + 2 Fill x with 2 : 2 + 2 = 4